Noluvpablo:

The work of a student to solve a set of equations is shown: Equation 1: m = 8 + 2n Equation 2: 3m = 4 + 4n Step 1: −3(m) = −3(8 + 2n) [Equation 1 is multiplied by −3.] 3m = 4 + 4n [Equation 2] Step 2: −3m = −24 − 6n [Equation 1 in Step 1 is simplified.] 3m = 4 + 4n [Equation 2] Step 3: −3m + 3m = −24 − 6n + 4n [Equations in Step 2 are added.] Step 4: 0 = −24 − 2n Step 5: n = −12 In which step did the student first make an error? Step 4 Step 3 Step 2 Step 1

snowflake0531:

could you put a screenshot, it makes it easier to read

jhonyy9:

Step 2: −3m = −24 − 6n [Equation 1 in Step 1 is simplified.] 3m = 4 + 4n [Equation 2] Step 3: −3m + 3m = −24 − 6n + 4n [Equations in Step 2 are added.] are added the left part is right -3m +3m = -24 -6n + from here missed the 4 added just 4n hope you get it

jhonyy9:

@AZ any idea pls - ty